### v1

@@ Abstract
Understanding feature-outcome associations in high-dimensional data remains challenging when relationships vary across subpopulations, yet standard methods assuming global associations miss context-dependent patterns, reducing statistical power and interpretability. We develop a geometric decomposition framework offering two strategies for partitioning inference problems into regional analyses on data-derived Riemannian graphs. Gradient flow decomposition uses path-monotonicity-validated discrete Morse theory to partition samples into basins where outcomes exhibit monotonic behavior. Co-monotonicity decomposition leverages association structure: vertex-level coefficients measuring directional concordance between outcome and features, or between feature pairs, define embeddings of samples into association space. These embeddings induce Riemannian k-NN graphs on which biclustering identifies co-monotonicity cells (coherent regions) and feature modules. This extends naturally to multi-modal integration across multiple feature sets. Both strategies apply independently or jointly, with Bayesian posterior sampling providing credible intervals.
@@ 2.1 From Point Clouds to Weighted Graphs
The k-nearest neighbor graph provides a more geometric perspective. We connect each point to its $k$ nearest neighbors, creating a graph $G=(V,E)$ where vertices correspond to observations and edges encode proximity relationships. For our two-cluster example with $k=4$, edges connect points within each cluster but not across clusters (assuming sufficient separation). The graph structure reveals the discrete geometry: the two clusters appear as connected components, and any function defined on vertices can diffuse within components but not between them.
However, unweighted graphs lose important information. In regions where points lie densely, small Euclidean distances separate neighbors, while in sparse regions, large distances occur. A single long edge might connect distant points that happen to be mutual nearest neighbors, creating an artificial bridge that permits gradient flow to jump between geometrically distant regions. We require vertex and edge weights that encode both local density and geometric scale.
@@ 2.2 Riemannian Structure Through Mass Assignment
Let $N_{k}(x_{i})$ denote the closed k-nearest neighbor ball of observation $x_{i}$, and let $\mathcal{U}=\{N_{k}(x_{1}),\ldots,N_{k}(x_{n})\}$ be the covering of the dataset by these neighborhoods. The nerve complex of this covering has vertices corresponding to observations, with simplices recording multi-way neighborhood intersections. We focus on the 1-skeleton (vertices and edges), which forms a graph where an edge $[i,j]$ exists whenever $N_{k}(x_{i})\cap N_{k}(x_{j})\neq\emptyset$.
Definition 1 (Vertex Mass).
The vertex mass $m_{0}(i)$ represents the measure of the neighborhood $N_{k}(x_{i})$. In practice, we use density-surrogate weights
@@
$m_{1}([i,j])=\sum_{\ell:x_{\ell}\in N_{k}(x_{i})\cap N_{k}(x_{j})}m_{0}(\ell).$ (2)
More generally, for two edges $e_{ij}=[i,j]$ and $e_{is}=[i,s]$ sharing vertex $i$, their inner product equals the triple intersection mass:
$\langle e_{ij},e_{is}\rangle=\sum_{\ell:x_{\ell}\in N_{k}(x_{i})\cap N_{k}(x_{j})\cap N_{k}(x_{s})}m_{0}(\ell).$ (3)
These inner products encode geometric relationships beyond simple edge weights. Two edges sharing a vertex are orthogonal if their corresponding neighborhoods intersect only at their common vertex, while edges with large triple intersection mass form acute angles. The collection of all edge inner products assembles into the edge mass matrix $M_{1}$, with diagonal entries $M_{1}(e,e)=m_{1}(e)$ and off-diagonal entries $M_{1}(e_{ij},e_{is})=\langle e_{ij},e_{is}\rangle$.
@@ 2.3 The Graph Laplacian and Diffusion
The Riemannian structure induces differential operators through the Hodge decomposition. We begin with the boundary operator $B_{1}:C_{1}\to C_{0}$, which maps edge chains to vertex chains according to
$(B_{1}\alpha)(i)=\sum_{j\in N(i)}(\alpha([j,i])-\alpha([i,j])),$ (4)
where $N(i)$ denotes the neighborhood of vertex $i$ and $[i,j]$ denotes the oriented edge from $i$ to $j$. The adjoint operator $B_{1}^{*}:C_{0}\to C_{1}$ with respect to the inner products defined by $M_{0}$ and $M_{1}$ satisfies
$\langle B_{1}\alpha,f\rangle_{M_{0}}=\langle\alpha,B_{1}^{*}f\rangle_{M_{1}}$ (5)
for all edge chains $\alpha$ and vertex chains $f$. This determines
$B_{1}^{*}=M_{1}^{-1}B_{1}^{T}M_{0}.$ (6)
Definition 3 (Hodge Laplacian).
The Hodge Laplacian on vertex chains (the graph Laplacian) is defined as
$L_{0}=B_{1}^{*}B_{1}=M_{0}^{-1}B_{1}^{T}M_{1}^{-1}B_{1}.$ (7)
This operator governs diffusion on the weighted graph. For computational purposes, we work with the symmetrized normalized Laplacian
$L_{\text{norm}}=M_{0}^{-1/2}L_{0}M_{0}^{-1/2}=M_{0}^{-1/2}B_{1}^{T}M_{1}^{-1}B_{1}M_{0}^{-1/2},$ (8)
which has eigenvalues in $[0,2]$ and admits spectral decomposition through standard symmetric eigensolvers.
The heat equation $\partial\rho/\partial t=-L_{0}\rho$ describes how vertex masses evolve under diffusion. The solution $\rho(t)=\exp(-tL_{0})\rho(0)$ applies the heat kernel to the initial distribution. For small $t$, diffusion smooths local irregularities while preserving global structure; for large $t$, all mass flows toward the stationary distribution. The Riemannian structure determines diffusion rates: edges with large mass (small Riemannian length) facilitate rapid exchange, while edges with small mass (large Riemannian length) impede flow.
@@ 2.4 Functions on Graphs and the Discrete Gradient
Definition 4 (Discrete Gradient).
For a function $f:V\to\mathbb{R}$, the gradient is the linear operator $\nabla f:E\to\mathbb{R}$ defined on oriented edges. For an edge $e=[i,j]$ directed from vertex $i$ to vertex $j$, we define
$(\nabla f)(e)=\frac{m_{0}(j)f(j)-m_{0}(i)f(i)}{m_{1}(e)},$ (9)
@@
This formula has intuitive interpretation. In regions where both vertices have equal mass (uniform density), the gradient reduces to the function difference divided by edge mass, analogous to a directional derivative where edge mass acts as effective distance. In non-uniform regions, vertices with larger mass contribute more strongly, reflecting that these vertices represent denser neighborhoods where the function value has greater statistical weight.
The gradient satisfies a fundamental variational property. Among all edge-valued functions $\phi$ with specified values at vertices, the gradient minimizes the energy
$\mathcal{E}(\phi)=\sum_{e\in E}m_{1}(e)|\phi(e)|^{2},$ (10)
subject to the boundary constraint $(B_{1}\phi)(i)=\sum_{j\in N(i)}m_{0}(j)(\phi([i,j])-\phi([j,i]))=0$ for all interior vertices. This characterizes the gradient as the minimizer of weighted squared variation, balancing smoothness against fidelity to vertex values.
2.5 Critical Points and the Discrete Morse Complex
A vertex $v$ is a local extremum if the function increases or decreases strictly when moving to any neighbor. Formally, $v$ is a local minimum if $f(v)<f(u)$ for all $u\in N(v)$, and a local maximum if $f(v)>f(u)$ for all $u\in N(v)$. These definitions extend classical Morse theory to the discrete setting.
Definition 5 (Critical Points).
The set of critical points consists of all local minima, local maxima, and (in certain formulations) saddle points where the Hessian has both positive and negative eigenvalues. For functions on graphs, we identify local extrema directly from the vertex function values through neighborhood comparisons.
Critical points organize the qualitative structure of the function. Between any local minimum $m$ and local maximum $M$, there exists a connecting path along which the function increases monotonically. These monotone paths define the gradient flow structure.
Definition 6 (Gradient Flow Trajectories).
Starting from a vertex $v$, the ascending gradient trajectory follows edges of steepest ascent: at each step, we move to the neighbor $u$ with maximal function value $f(u)>f(v)$, continuing until reaching a local maximum where no higher neighbor exists. Similarly, the descending trajectory follows steepest descent to a local minimum.
We denote the terminus of the ascending trajectory from $v$ as $\pi_{\uparrow}(v)$, and the terminus of the descending trajectory as $\pi_{\downarrow}(v)$. These maps partition the vertex set into basins of attraction.
Definition 7 (Basins of Attraction).
For a local minimum $m$, the ascending basin is
$A(m)=\{v\in V:\pi_{\uparrow}(v)=m\},$ (11)
containing all vertices whose ascending flow terminates at $m$. For a local maximum $M$, the descending basin is
$D(M)=\{v\in V:\pi_{\downarrow}(v)=M\},$ (12)
containing all vertices whose descending flow terminates at $M$.
The basins satisfy $\bigcup_{m}A(m)=V$ and $\bigcup_{M}D(M)=V$, providing complete partitions of the vertex set. The intersections $C(m,M)=A(m)\cap D(M)$ define gradient flow cells, which are regions where both ascending and descending flows have unique termini.
Theorem 8 (Monotonicity in Gradient Flow Cells).
Within a gradient flow cell $C(m,M)$, the function $f$ is monotone along all paths connecting the minimum $m$ to the maximum $M$. Specifically, any path $m=v_{0},v_{1},\ldots,v_{k}=M$ contained in $C(m,M)$ satisfies $f(v_{0})<f(v_{1})<\cdots<f(v_{k})$.
The theorem follows from the definitions: within $C(m,M)$, ascending flow from any vertex moves toward $M$, so function values increase along ascending paths; descending flow moves toward $m$, so function values decrease along descending paths. Any monotone path from $m$ to $M$ must ascend throughout, ensuring strict increase.
This monotonicity property provides the foundation for regional inference. Within each cell, the outcome exhibits uniform qualitative behavior (monotone increase from minimum to maximum), suggesting that feature associations should also exhibit spatial homogeneity within cells but may differ across cells.
2.6 The Long Edge Problem and Path Monotonicity
Real data graphs frequently contain long edges connecting geometrically distant points that happen to be mutual nearest neighbors. These edges create artifacts in gradient flow computation: a trajectory may follow a single long edge to jump between basins, violating the assumption that flow follows the intrinsic geometry of the data manifold.
We address this through path monotonicity validation. For an edge $e=[i,j]$ with $f(i)<f(j)$, we compute the shortest alternative path $P(i,j)$ that avoids $e$. If $f$ increases strictly along $P(i,j)$, the edge respects the underlying monotonicity and participates in gradient flow. If $f$ oscillates along $P(i,j)$, the edge creates an artificial shortcut and we exclude it from flow computation, requiring trajectories to follow longer but geometrically faithful paths.
Formally, we define the validated edge set
$E_{\text{val}}=\{[i,j]\in E:f\text{ is monotone along }P(i,j)\}$ (13)
and compute gradient flow using only edges in $E_{\text{val}}$. This removes spurious connections while preserving genuine geometric structure.
The path monotonicity criterion has theoretical justification. If the function $f$ truly varies smoothly along the data manifold, then any edge connecting points with $f(i)<f(j)$ should lie in a region where $f$ increases throughout. The existence of a non-monotone alternative path indicates that the direct edge cuts across level sets rather than following the manifold geometry, suggesting that it represents noise in the k-nearest neighbor construction rather than intrinsic structure.
@@ 2.7 Spectral Filtering and Conditional Expectation Estimation
2.8 Summary of the Geometric Framework
We have constructed a complete framework for representing and analyzing functions on data-derived graphs. The Riemannian structure $(M_{0},M_{1})$ encodes local density and geometric relationships through vertex and edge masses. The graph Laplacian $L_{0}$ governs diffusion and enables spectral filtering for robust estimation. The discrete gradient provides a natural notion of directional derivative, leading to gradient flow trajectories and basin decomposition. Path monotonicity validation removes artifactual long edges, ensuring that gradient flow follows the intrinsic data geometry.
This geometric perspective transforms the statistical inference problem. Rather than estimating $E[Y|X]$ in the ambient Euclidean space $\mathbb{R}^{d}$, we estimate a function on the graph $G(X)$ and partition the graph into regions of monotonic behavior through gradient flow analysis. The partition emerges from the geometry rather than being imposed a priori, and the Riemannian structure ensures that inference respects the intrinsic dimensionality and density variation of the data.
We now proceed to develop the complete inference framework: identification and filtering of spurious extrema, clustering of basins to manage over-segmentation, and introduction of co-monotonicity measures for association analysis within gradient flow cells.
@@ 3.1 The Challenge of Long Edges
Consider constructing a k-nearest neighbor graph from a sample drawn from a low-dimensional manifold embedded in high-dimensional space. The standard construction connects each point to its k nearest neighbors by Euclidean distance in the ambient space, creating edges that respect local density but potentially violate global geometry. When the manifold curves or has boundaries, points that are close in Euclidean distance may lie far apart along the manifold. A point near the edge of a curved region might have as its k-th nearest neighbor a point on the opposite side of a valley, leading to an edge that shortcuts across the manifold rather than following its intrinsic geometry.
These long edges are necessary for graph connectivity. Without them, the graph fragments into components corresponding to dense regions, preventing paths between different parts of the sample. Yet these same edges create artifacts in gradient flow computation. Consider a vertex $v$ in a descending basin of a local minimum $m$, where the smoothed function $\hat{y}$ decreases monotonically along all paths through neighboring vertices toward $m$. If $v$ has a long edge to a vertex $u$ in a different basin where $\hat{y}(u)<\hat{y}(v)$, the naive gradient rule (follow the neighbor with minimal function value) directs the trajectory to jump across basins, violating the fundamental property that gradient flow should follow monotone paths along the underlying geometry.
We illustrate this phenomenon through a concrete example. Let $M$ be a manifold in $\mathbb{R}^{2}$ consisting of two crescents separated by a narrow gap, with a function $f:M\to\mathbb{R}$ having a local minimum in each crescent. Dense sampling near the crescent tips yields k-nearest neighbor graphs with edges connecting the two components. For a vertex $v$ near one crescent tip, its nearest neighbors in the ambient Euclidean metric include points across the gap in the other crescent, despite the manifold distance being substantially larger. When $v$ lies in the basin of minimum $m_{1}$ in the first crescent, but has a neighbor $u$ in the basin of minimum $m_{2}$ in the second crescent with $\hat{y}(u)<\hat{y}(v)$, the naive gradient trajectory jumps from basin to basin, producing what we term basin jumping artifacts.
@@ 3.2 Inadequacy of Existing Approaches
One might attempt to resolve this issue through the Hodge Laplacian gradient, which incorporates global geometric information through the mass matrices. Recall that the discrete gradient of a function $y:V\to\mathbb{R}$ can be computed as a 1-chain (function on edges) via
$(\nabla y)_{e}=[M_{1}^{-1}B_{1}^{T}M_{0}y]_{e},$ (16)
where $B_{1}:\mathbb{R}^{E}\to\mathbb{R}^{V}$ is the coboundary operator, $M_{0}$ is the vertex mass matrix, and $M_{1}$ is the edge mass matrix encoding the Riemannian metric. This gradient assigns to each edge a value that reflects not merely the local function difference but the edge’s role in the global geometric structure.
The edge mass $M_{1}[e,e]$ for an edge $e=[i,j]$ quantifies the overlap between the k-neighborhoods of vertices $i$ and $j$, with larger mass indicating stronger geometric connection. Long edges connecting distant points have small mass (small neighborhood overlap), and the inverse $M_{1}^{-1}$ amplifies their contribution, potentially destabilizing the gradient computation. While this approach implicitly downweights long edges through their small mass values, it provides only an indirect solution. Computing the Hodge gradient requires solving a linear system involving $M_{1}^{-1}$, which can be ill-conditioned when edge masses vary over several orders of magnitude, a common occurrence in real data where sampling density fluctuates. Moreover, the spectral properties of $M_{1}$ do not directly enforce the path monotonicity that characterizes true gradient trajectories.
Alternatively, one might restrict gradient flow computation to edges shorter than a quantile threshold $\tau_{q}$, where $\tau_{q}$ is the $q$-quantile of the edge length distribution for some $q\in[0.5,1]$. This heuristic removes the most problematic edges but discards potentially valid geometric information. A long edge might legitimately connect vertices that lie on a genuine gradient trajectory if the underlying function varies smoothly between them. Uniform length thresholding cannot distinguish between geometrically faithful long edges and artifactual ones, leading to unnecessary loss of connectivity and potential fragmentation of basins.

### v2

@@ Abstract
Understanding feature-outcome associations in high-dimensional data remains challenging when relationships vary across subpopulations, yet standard methods assuming global associations miss context-dependent patterns, reducing statistical power and interpretability. We develop a geometric decomposition framework offering two strategies for partitioning inference problems into regional analyses on data-derived Riemannian graphs. Gradient flow decomposition uses path-monotonicity-validated discrete Morse theory to partition samples into gradient flow cells where outcomes exhibit monotonic behavior. Co-monotonicity decomposition leverages association structure: vertex-level coefficients measuring directional concordance between outcome and features, or between feature pairs, define embeddings of samples into association space. These embeddings induce Riemannian k-NN graphs on which biclustering identifies co-monotonicity cells (coherent regions) and feature modules. This extends naturally to multi-modal integration across multiple feature sets. Both strategies apply independently or jointly, with Bayesian posterior sampling providing credible intervals.
@@ 2.1 From Point Clouds to Weighted Graphs
The intersection k-nearest neighbor (ikNN) graph $G_{k}(X)$ associated with observations $X\subset\mathbb{R}^{d}$ provides a more geometric perspective. The vertices of $G_{k}(X)$ are all points $X$ and two vertices $x_{i},x_{j}$ are connected by an edge if and only if $\hat{N}_{k}(x_{i})\cap\hat{N}_{k}(x_{j})\neq\emptyset$, where $\hat{N}_{k}(x_{i})$ is the closed k-nearest neighbor ball of $x_{i}$, consisting of all k-nearest neighbors of $x_{i}$ and the point $x_{i}$. For our two-cluster example with $k=4$, edges connect points within each cluster but not across clusters (assuming sufficient separation). The graph structure reveals the discrete geometry: the two clusters appear as connected components, and any function defined on vertices can diffuse within components but not between them.
However, unweighted graphs lose important information. In regions where points lie densely, small Euclidean distances separate neighbors, while in sparse regions, large distances occur. A single long edge might connect distant points creating an artificial bridge that permits gradient flow to jump between geometrically distant regions. We require vertex and edge weights that encode both local density and geometric scale.
@@ 2.2 Riemannian Structure Through Mass Assignment
Let $\mathcal{U}=\{\hat{N}_{k}(x_{1}),\ldots,\hat{N}_{k}(x_{n})\}$ be the kNN balls covering of $X=\{x_{1},x_{2},\ldots,x_{n}\}$. The nerve complex of this covering has vertices corresponding to observations, with simplices recording multi-way neighborhood intersections. We focus on the 1-skeleton (vertices and edges), which forms a graph where an edge $[i,j]$ exists whenever $\hat{N}_{k}(x_{i})\cap\hat{N}_{k}(x_{j})\neq\emptyset$. Thus, the intersection k-nearest neighbor graph is the 1-skeleton of the nerve complex of the k-nearest neighbor covering.
The vertex mass $m_{0}(i)$ represents the measure of the neighborhood $\hat{N}_{k}(x_{i})$. In practice, we use density-surrogate weights
@@
$m_{1}([i,j])=\sum_{\ell:x_{\ell}\in\hat{N}_{k}(x_{i})\cap\hat{N}_{k}(x_{j})}m_{0}(\ell).$ (2)
For two edges $e_{ij}=[i,j]$ and $e_{is}=[i,s]$ sharing vertex $i$, their inner product equals the triple intersection mass:
$\langle e_{ij},e_{is}\rangle=\sum_{\ell:x_{\ell}\in\hat{N}_{k}(x_{i})\cap\hat{N}_{k}(x_{j})\cap\hat{N}_{k}(x_{s})}m_{0}(\ell).$ (3)
Thus, $e_{ij},e_{is}$ are orthogonal if and only if
$\hat{N}_{k}(x_{i})\cap\hat{N}_{k}(x_{j})\cap\hat{N}_{k}(x_{s})=\emptyset$
assuming $m_{0}(x_{i})>0$ for all $x_{i}$.
These inner products encode geometric relationships beyond simple edge weights. Two edges sharing a vertex are orthogonal if their corresponding neighborhoods intersect trivially, while edges with large triple intersection mass form acute angles. The collection of all edge inner products assembles into the edge mass matrix $M_{1}$, with diagonal entries $M_{1}(e,e)=m_{1}(e)$ and off-diagonal entries $M_{1}(e_{ij},e_{is})=\langle e_{ij},e_{is}\rangle$.
@@ 2.3 The Graph Laplacian and Diffusion
The Riemannian structure determines a graph Laplacian that governs diffusion on the weighted graph. We begin with the boundary operator $B_{1}:C_{1}\to C_{0}$, which maps edge chains to vertex chains according to
$(B_{1}\alpha)(i)=\sum_{j\in N(i)}(\alpha([j,i])-\alpha([i,j])),$
where $N(i)$ denotes the neighborhood of vertex $i$ and $[i,j]$ denotes the oriented edge from $i$ to $j$. This operator encodes the graph’s combinatorial structure through the vertex-edge incidence relations.
We define the graph Laplacian using edge conductances derived from the edge mass matrix. For each edge $e$, we set the conductance $c_{e}=m_{1}(e)$, using the edge mass directly. Since the base measure $m_{0}(i)=(\varepsilon+d_{k}(x_{i}))^{-\alpha}$ already inverts local scales, edges in dense regions naturally receive large conductances while edges in sparse regions receive small conductances. Assembling these into a diagonal matrix $C=\text{diag}(c_{1},\ldots,c_{m})$, we construct the unnormalized Laplacian
$L_{\text{div}}=B_{1}CB_{1}^{T}.$
This operator is symmetric and positive semidefinite, with eigenvalues encoding the graph’s connectivity structure.
For computational purposes, we work with the symmetrized normalized Laplacian
$L_{\text{norm}}=M_{0}^{-1/2}L_{\text{div}}M_{0}^{-1/2},$
which has eigenvalues in $[0,2]$ and admits spectral decomposition through standard symmetric eigensolvers. The normalization by vertex masses balances the influence of vertices with different local densities, preventing high-degree vertices from dominating the diffusion process.
The heat equation $\partial\rho/\partial t=-L_{\text{div}}\rho$ describes how vertex masses evolve under diffusion. The solution $\rho(t)=\exp(-tL_{\text{div}})\rho(0)$ applies the heat kernel to the initial distribution. For small $t$, diffusion smooths local irregularities while preserving global structure; for large $t$, all mass flows toward the stationary distribution. The Riemannian structure determines diffusion rates: edges with large mass (small Riemannian length) facilitate rapid exchange, while edges with small mass (large Riemannian length) impede flow.
The edge mass matrix $M_{1}$ contains not only diagonal entries $m_{1}(e)$ but also off-diagonal entries encoding inner products between edges sharing vertices through triple neighborhood intersections. In the Laplacian construction above, we use only the diagonal structure for computational efficiency. The full Riemannian geometry, including these off-diagonal terms, is incorporated in the gradient computation used for basin analysis (Section 3), where it plays a crucial role in determining flow directions. Extending the Laplacian construction to utilize the complete $M_{1}$ structure remains an important direction for future work.
@@ 2.4 Functions on Graphs and the Discrete Gradient
The gradient operator $\nabla f:E\to\mathbb{R}$ assigns to each oriented edge a value encoding the directional rate of change of $f$. The Riemannian structure determines this assignment through the adjoint of the boundary operator. Formally, the gradient is defined by the relation
$\langle B_{1}\phi,f\rangle_{M_{0}}=\langle\phi,\nabla f\rangle_{M_{1}}$
for all edge functions $\phi$ and vertex functions $f$, where the left side uses the vertex inner product weighted by $M_{0}$ and the right side uses the edge inner product weighted by $M_{1}$. This yields $\nabla f=B_{1}^{*}(f)$, where $B_{1}^{*}:C_{0}\to C_{1}$ denotes the adjoint operator.
In the diagonal case, where edges are orthogonal in the Riemannian metric (meaning $M_{1}$ is diagonal with $M_{1}(e_{ij},e_{is})=0$ for distinct edges sharing a vertex), the gradient has an explicit formula. For an edge $e=[i,j]$ directed from vertex $i$ to vertex $j$, we have
$(\nabla f)(e)=\frac{m_{0}(j)f(j)-m_{0}(i)f(i)}{m_{1}(e)},$
@@
This formula admits a natural interpretation. In regions where both vertices have equal mass (uniform density), the gradient reduces to the function difference divided by edge mass, analogous to a directional derivative where edge mass acts as effective distance. In non-uniform regions, vertices with larger mass contribute more strongly, reflecting that these vertices represent denser neighborhoods where the function value carries greater statistical weight. The division by edge mass ensures that the gradient measures rate of change per unit geometric distance rather than absolute difference.
When the full non-diagonal structure of $M_{1}$ is employed (incorporating off-diagonal entries from triple neighborhood intersections), computing $\nabla f=B_{1}^{*}(f)$ requires solving the linear system $M_{1}x=B_{1}^{T}M_{0}f$, which we address through iterative methods in the basin analysis of Section 3.
@@ 3.1 The Challenge of Long Edges
Consider constructing a k-nearest neighbor graph from a sample drawn from a low-dimensional manifold embedded in high-dimensional space. When the manifold curves or has boundaries, points that are close in Euclidean distance may lie far apart along the manifold. A point near the edge of a curved region might have as its k-th nearest neighbor a point on the opposite side of a valley, leading to an edge that shortcuts across the manifold rather than following its intrinsic geometry.
These long edges are necessary for graph connectivity. Without them, the graph fragments into components corresponding to dense regions, preventing paths between different parts of the sample. Yet these same edges can create artifacts in gradient flow computation. Consider a vertex $v$ in a descending basin of a local minimum $m$, where the smoothed function $\hat{y}$ decreases monotonically along all paths through neighboring vertices toward $m$. If $v$ has a long edge to a vertex $u$ in a different basin where $\hat{y}(u)<\hat{y}(v)$, the naive gradient rule (follow the neighbor with minimal function value) directs the trajectory to jump across basins, violating the fundamental property that gradient flow should follow monotone paths along the underlying geometry.
@@ 3.2 Inadequacy of Existing Approaches
One might attempt to resolve this issue through more sophisticated gradient computation. For basin analysis, we can compute the discrete gradient incorporating the full Riemannian structure via
$(\nabla y)_{e}=[M_{1}^{-1}B_{1}^{T}M_{0}y]_{e},$
where $B_{1}:\mathbb{R}^{E}\to\mathbb{R}^{V}$ is the boundary operator, $M_{0}$ is the vertex mass matrix, and $M_{1}$ is the full edge mass matrix including off-diagonal entries encoding geometric relationships between edges sharing vertices (see Appendix). This gradient accounts for global geometric structure beyond local function differences.
However, even this sophisticated gradient computation does not resolve the long edge problem. The gradient formula incorporates edge masses that quantify neighborhood overlaps, but **the fundamental issue remains**: a long edge connecting vertices in different basins can still produce a gradient that directs flow across basin boundaries, even when the edge has appropriately small mass. The problem is not that the gradient is computed incorrectly, but that the graph topology itself contains edges that violate the manifold geometry. No local gradient computation can distinguish between a long edge that shortcuts across a valley (problematic) versus one that legitimately connects distant points along a genuine gradient path (acceptable).
Alternatively, one might restrict gradient flow to edges shorter than a quantile threshold $\tau_{q}$, where $\tau_{q}$ is the $q$-quantile of the edge length distribution for some $q\in[0.5,1]$. This heuristic removes the most problematic edges but discards potentially valid geometric information. A long edge might legitimately connect vertices that lie on a genuine gradient trajectory if the underlying function varies smoothly between them. Uniform length thresholding cannot distinguish between geometrically faithful long edges and artifactual ones, leading to unnecessary loss of connectivity and potential fragmentation of basins.
What we need instead is a criterion that validates each edge based on the actual function behavior along paths, not merely on edge length or local gradient values. This motivates our path monotonicity approach.